OpenStudy (anonymous):

Carl takes a random sample of 100 students in his school and finds that 72% of the sample prefers pizza over burgers. There are 1,500 students in the school. Based on the sample proportion, how many students in the school would be expected to prefer pizza over burgers?

OpenStudy (anonymous):

choices are A. 72 B. 100 C. 1,080 D. 1,500

OpenStudy (anonymous):

will fan and medal

OpenStudy (ashes.boom):

\[\frac{ 72 }{ 100 } = \frac{ x }{ 1500 }\] cross multiply: 72*1500 = 100x 108000 = 100x divide both sides by 100 to get x 108000/100 = 100x / 100 1080 = x

OpenStudy (anonymous):

so 1,080 is my answer

OpenStudy (ashes.boom):

yep

OpenStudy (anonymous):

thx i medaled and fanned you

OpenStudy (ashes.boom):

ur welcome and thanks :)